# In-Context Imitation Learning via Next-Token Prediction

### v1

@@ 4.2 Model Architecture
To facilitate in-context learning in a robotics setting, the model should have a sufficiently long context window to support prompting by providing robot trajectories as demonstrations. We construct the ICRT model with three parts: a pre-trained vision encoder, a series of projectors for each input modality, and a causal transformer backbone (Figure 2).
@@
Vision Encoder The model processes multi-view image observations through a pre-trained vision transformer. However, most visual pre-trained networks are trained on ImageNet or human videos [27, 51, 52, 24], which exhibit a significant domain gap when compared to typical images from robot datasets, where the images frequently include robots or grippers. To minimize the domain gap, we pre-train a vision transformer [53] (ViT-Base) on an equal mix of ImageNet [54] and Open X-Embodiment [40] data, using CrossMAE as an efficient pre-training method [55]. During the training of the ICRT model, we freeze the vision encoder for efficiency. The vision encoder outputs the entire feature map for each of the cameras and is then fed into the proprioception projector (Figure 2 left).
@@
Modality-Specific Projectors To project image observations, the robot’s proprioceptive state, and actions into a shared latent space for sequence modeling, we design modality-specific projectors. At each timestep, the model takes as input a token representing either the robot’s state or an action. To produce a single state token that captures fine-grained visual information and the proprioceptive state of the robot, we use attention pooling [56] between all visual tokens from a single camera’s observation and a proprioception embedding produced by a multi-layer perceptron (MLP). The resulting embeddings for each camera are concatenated to produce a single state token $f_{s}^{t}$ of dimension equal to the transformer latent dimension. Similar to proprioception, the action is embedded with an MLP into an action token $f_{a}^{t}$. This process produces a sequence of state and action tokens that are passed into the transformer.
@@
Transformer Model The encoded sequence of state and actions is passed into a Transformer model [57], following the design of Llama2 [12]. The transformer takes as input the sequence of state and action features $(f_{s}^{1},f_{a}^{1},\cdots,f_{s}^{t},f_{a}^{t})$ that are produced by the modality-specific projectors. We add MLP decoders to produce state and action outputs from the last layer of the transformer at the appropriate positions. We denote the transformer with the decoder heads as $g_{\theta}$. Therefore, the desired outputs are the shifted sequence of proprioceptive states and actions $(a^{1},s^{2},a^{2},\cdots,a^{t},s^{t+1})$. This naturally forms a next token prediction problem, as $g_{\theta}(f_{s}^{1})$ predicts $a^{1}$ and $g_{\theta}(f_{s}^{1},f_{a}^{1},\cdots,f_{s}^{n})$ predicts $a^{n+1}$. In practice, we find it beneficial to predict the next $h$ actions at each time step, and use temporal ensembling [2] to execute the final action.
Inspired by Octo [15] and vision transformers [53], we consider a randomly initialized Llama2 model of 12 layers with a latent dimension of 768, which we name Llama2-Base. In addition, multiple works have shown that multimodal inputs can be aligned to large-language models [34, 58, 59, 8, 60]. Multi-modal language model, Palm-E [10] has shown success in enhancing generalization when being directly incorporated into robotic control [8]. Therefore, we also investigate the effectiveness of using a large-language model for in-context robot learning by initializing the transformer with a pre-trained Llama2-7B. Due to the large domain gap between natural language and robot trajectories, a frozen language model may not be sufficient. Therefore, similar to prior work in multimodal alignment, we fine-tune the language model with LoRA [61], with a rank of 32. Due to compute resource limitations, we are unable to fully fine-tune the model.
Loss Function To provide more supervision signals so that the model can better respond to the trajectory “prompt” we provide at test time, we reference works in training multi-turn conversation chatbots [62, 34], where they only compute loss on the response generated by the chatbot, instead of the prompt. Recall that in Section 4.1, we randomly sampled the subsequence of the concatenated trajectories as the prompt trajectory. Analogously, we only compute action prediction with L1-loss for the actions after the prompt trajectories.
A key advantage of this framework is its use of the transformer’s sequential processing capability. Instead of reprocessing the entire sequence history for each model evaluation, as seen in previous works [15, 14, 7, 8], the model employs a key-value (KV) caching mechanism, as discussed in [12]. This mechanism stores previous outputs, allowing the model to compute only the outputs for the new token. This approach significantly reduces computational overhead, lowering the complexity from quadratic to linear relative to the sequence length.
5 Experiments
Figure 4: Physical experiments setup are shown on the left, showing the Franka Emika robot, the wrist and side camera and the objects used in training and evaluation. We consider 6 primitives and collect human demonstrations for training. We consider pick up and place and poke primitives for evaluation (dark green).
@@ 5 Experiments
In this section, we design an experimental setup to evaluate the in-context learning capabilities of the proposed models for continuous robot control and compare them against several baselines. Instead of focusing on the difficulty of learning a specific task primitive, we design the experiments to assess the policy’s ability to accomplish unseen tasks among all executable options from the provided prompt trajectories.
@@
Experiment Design We consider two action primitives: a pick-and-place primitive and a poking primitive. For each action primitive, we design six unseen tasks (as defined in Section 3), with three tasks evaluating in-domain object generalization and three evaluating on objects unseen during training (selected from radish, blue sponge, grey dog, and black dog, see Table 1 and Table 2).
@@
Each task is designed with five tiers of difficulty. In the pick-and-place primitive, the model is tasked with identifying which object to grasp and where to place it in a multi-object or multi-placement setting. The tiers are: 1) pick and place the object without any distractors, 2) with one distractor object, 3) with two distractor objects, 4) with three distractor objects, and 5) with one distractor placement position. For the poking primitive, the robot must close the gripper, poke the object, lift the end-effector, and open the gripper. The five tiers of difficulty involve the target object presented with 0-4 distractors in the scene.
@@
The pick-and-place primitive is evaluated by assigning a partial credit of 0.5 if the robot correctly picks up the object. A successful placement results in a total score of 1. The poking task is evaluated by whether the model pokes the correct object; if an incorrect object is poked, the trial is marked as a failure. The model is allowed retries within a time limit of 25 seconds (or 375 steps). Each tier of difficulty is performed once, and we report the average success rate per task, as well as the average success rate and standard deviation per action primitive across the six tasks.
@@
Algorithms The default ICRT model is a randomly initialized Llama2-Base model pretrained on DROID and fully fine-tuned on ICRT-MT. We evaluate the impact of model initialization and training datasets by introducing the following three variants: 1) ICRT-Llama2, a pre-trained Llama2-7B language model fine-tuned on ICRT-MT with LoRA; 2) ICRT (DROID), a randomly initialized Llama2-Base model trained only on the DROID dataset; and 3) ICRT (MT), a randomly initialized Llama2-Base model trained only on the ICRT-MT dataset.
@@
We consider 3 baseline algorithms. We train a goal-conditioned policy, where in each sample of the dataset, the goal observation and state pair are always prepended to the sequence, and in each sequence, there exists only one trajectory. This resembles the normal goal-conditioned imitation learning setup. Additionally, we finetune Octo [15], the state-of-the-art goal-image and language conditioned policy, and OpenVLA [14], the state-of-the-art language conditioned multi-task imitation learning algorithm. Octo is fine-tuned using their official fine-tuning recipe. We incorporate action chunking into OpenVLA by asking it to predict the next 16 actions, which performs better than vanilla OpenVLA which predicts only the next step. Both of these methods are representative of robot policies that use next-token prediction objectives.
@@
Prompt Generation For each task, we collect 3 demonstrations (with zero, one distractor object, a distractor placement for pick-and-place, or two distractor objects for poking) as the prompt in total before running the experiment. Please refer to the Appendix 8.1 Figure 5 for a visual example. During testing, a random demonstration is drawn as a prompt to assess the model’s ability to generalize to different prompts. It’s important to note that the environment setup during policy rollout differs from the prompts’ setup, ensuring that the evaluation measures the model’s understanding of task-relevant information from the prompt, rather than simply copying actions from it.
Pick and Place Pick Object Place Location Yellow Cube Black Bowl Yellow Cube Grey Bowl Blue Bear Pink Bowl Radish Grey Bowl Black Dog Pink Bowl Blue Sponge Silver Pot Average Success ($\pm$ Std.) Goal Conditioned 40% 30% 20% 40% 40% 30% 33.3% ($\pm$7.5%) Octo 10% 0% 10% 10% 0% 0% 5.0% ($\pm$5.0%) OpenVLA 0% 0% 0% 50% 20% 0% 11.7% ($\pm$18.6%) ICRT-Llama2 40% 40% 40% 60% 40% 40% 43.3% ($\pm$7.5%) ICRT (DROID) 0% 0% 0% 0% 0% 0% 0.0% ($\pm$0.0%) ICRT (MT) 90% 50% 80% 90% 60% 90% 76.7% ($\pm$16.0%) ICRT 60% 50% 80% 50% 60% 90% 65.0% ($\pm$15.0%)
Table 1: Pick up and place primitive performed with goal conditioned or by using one sequence as the prompt.
Poke Poke Object Radish Red Cube Grey Dog Black Cube Pink Bowl Blue Sponge Average Success ($\pm$ Std.) Goal Conditioned 0% 0% 0% 0% 40% 0% 6.7% ($\pm$14.9%) Octo 20% 0% 60% 0% 0% 0% 13.3% ($\pm$22.1%) OpenVLA 20% 0% 0% 0% 0% 0% 3.3% ($\pm$7.4%) ICRT-Llama2 60% 100% 80% 60% 60% 80% 73.3% ($\pm$14.9%) ICRT (DROID) 0% 0% 0% 0% 0% 0% 0.0% ($\pm$0.0%) ICRT (MT) 100% 100% 40% 60% 60% 60% 70.0% ($\pm$22.4%) ICRT 100% 100% 80% 80% 100% 100% 93.3% ($\pm$9.4%)
Table 2: Poking primitive performed with goal conditioned or by using one sequence as the prompt.
@@
Results We present the results in Table 1 and Table 2. For the pick-and-place primitive, we observe that the goal-conditioned policy generally succeeds in identifying the correct object to grasp when no distractor objects are present. However, its performance degrades significantly as the number of distractors increases. When the goal image only specifies the task but not the specific way to achieve it in the current environment, goal-conditioned policies often fail to execute the task effectively.
@@
The results suggest that ICRT outperforms the goal-conditioned policy in identifying the correct object to pick up and the appropriate placement location. The poking task presents a significant challenge for the goal or language-conditioned policies, as the goal position often closely resembles the start configuration. However, after conditioning on the prompt trajectory, ICRT is able to correctly identify the task as poking, and the results indicate that it consistently reaches the correct target object while ignoring distractors. Despite this, we do observe some failure modes with ICRT, such as missing the grasp of the target object, grasping the wrong object, or placing objects in incorrect locations. Specifically, when a distractor object shares the same color but has a different shape, the model struggles to accurately determine which object to grasp. This implies that additional fine-tuning of the vision encoder might be required to enhance model performance, a conclusion also reached by OpenVLA [14].
6 Ablations
@@
In this section, we provide additional experiments presented Table 1 and Table 2 that ablate on a few core design choices. We provide more ablation studies in Section 8.2.
6.1 Model Initialization
@@
We conducted ablation studies to examine the impact of using a pretrained Llama2 on language data and fine-tune it for robot sensorimotor sequence modeling. The results, presented in Table 1 and Table 2, show that although ICRT-Llama2-7B achieves a lower training loss, its performance is worse compared to its smaller counterparts. This discrepancy may be attributed to a lower inference frequency of ICRT-Llama2. We suggest that future work should focus on optimizing the inference speed of ICRT-Llama2.
6.2 Training Dataset
@@
We find that training on the DROID subset (see Section 4.1) is insufficient for successfully completing any of the test tasks; the policy (ICRT (DROID)) shows no progress across all tasks. This suggests that although the DROID subset may offer greater visual diversity, the unique structure of ICRT-MT—where multiple tasks are performed from the same initial observation—is particularly beneficial in developing the in-context learning capabilities of a next-token prediction robot model.
@@
ICRT (MT) shows similar performance to ICRT that is pretrained on DROID, especially for the pick-up and place primitive, even surpassing ICRT on the put radish in grey bowl task. However, ICRT (MT) does not perform as well on the poking primitive. The results suggest that it may be beneficial to pre-train the autoregressive model on a large dataset, as a diverse dataset may help the transformer to perform better alignment between visual features and control.
6.3 No Prompt Loss
@@
Following the design of many multi-turn conversation large language models or vision language model fine-tuning works [34, 62, 63, 64], we do not calculate the loss for the predicted action in the prompt trajectories but only do so on the predictions after the prompt trajectories. We mark the model that calculates loss on the prompt as ICRT +Prompt Loss and the default model as ICRT. The results are shown in Table 3 and Table 4. We find that by letting the model only predict the trajectories after the designated prompt trajectories, the model’s performance improves significantly. We hypothesize that in the situation where there is a loss on the prompt trajectories, the model is forced to do un-conditional generation based on the current environment observation, especially when there are multiple possible tasks available. This may cause the model to stop paying attention to the prompt.
Pick and Place
Pick Object
Place Location
Yellow Cube
Black Bowl
Yellow Cube
Grey Bowl
Blue Bear
Pink Bowl
Radish
Grey Bowl
Black Dog
Pink Bowl
Blue Sponge
Silver Pot
ICRT +Prompt Loss 20% 10% 20% 40% 30% 10%
ICRT 60% 50% 80% 50% 60% 90%
Table 3: Ablation on not applying loss on the prompt trajectories for pick and place tasks.
Poke
Poke Object Radish Red Cube Grey Dog Black Cube Pink Bowl Blue Sponge
ICRT +Prompt Loss 0% 20% 20% 80% 0% 20%
ICRT 100% 100% 80% 80% 100% 100%
Table 4: Ablation on not applying loss on the prompt trajectories for poking tasks.
7 Limitations and Conclusion

### v2

@@ IV-B Model Architecture
We construct the ICRT model with three parts: a pre-trained vision encoder, a series of projectors for each input modality, and a causal transformer backbone (Figure 3).
@@
Vision Encoder The model processes multi-view image observations through a pre-trained vision transformer. However, most visual pre-trained networks are trained on ImageNet or human videos [radosavovic2022, vc2023, chen2024sugar, nair2022], which exhibit a significant domain gap when compared to typical images from robot datasets, where the images frequently include robots or grippers. To minimize the domain gap, we pre-train a vision transformer [Dosovitskiy2020] (ViT-Base) on an equal mix of ImageNet [Deng2009] and Open X-Embodiment [2024rtx] data, using CrossMAE as an efficient pre-training method [fu2024rethinking]. During the training of the ICRT model, we freeze the vision encoder for efficiency. The vision encoder outputs the entire feature map for each of the cameras and is then fed into the proprioception projector (Figure 3 left).
@@
Modality-Specific Projectors To project image observations, the robot’s proprioceptive state, and actions into a shared latent space for sequence modeling, we design modality-specific projectors. At each timestep, the model takes as input a token representing either an observation or an action. To produce a single state token that captures fine-grained visual information and the proprioceptive state of the robot, we use attention pooling [lee2019set] between all visual tokens from a single camera’s observation and a proprioception embedding produced by a multi-layer perceptron (MLP). The resulting embeddings for each camera are concatenated to produce a single state token $f_{s}^{t}$ of dimension equal to the transformer latent dimension. Similar to proprioception, the action is embedded with an MLP into an action token $f_{a}^{t}$. This process produces a sequence of state and action tokens that are passed into the transformer.
Transformer Model The encoded sequence of state and actions is passed into a Transformer model [Vaswani2017], following the design of Llama2 [touvron2023llama]. The transformer takes as input the sequence of state and action features $(f_{s}^{1},f_{a}^{1},\cdots,f_{s}^{t},f_{a}^{t})$ that are produced by the modality-specific projectors. We add MLP decoders to produce state and action outputs from the last layer of the transformer at the appropriate positions. We denote the transformer with the decoder heads as $g_{\theta}$. Therefore, the desired outputs are the shifted sequence of proprioceptive states and actions $(a^{1},s^{2},a^{2},\cdots,a^{t},s^{t+1})$. This naturally forms a next token prediction problem, as $g_{\theta}(f_{s}^{1})$ predicts $a^{1}$ and $g_{\theta}(f_{s}^{1},f_{a}^{1},\cdots,f_{s}^{n})$ predicts $a^{n+1}$. In practice, we find it beneficial to predict the next $h$ actions at each time step, and use temporal ensembling [zhao2023learning] to execute the final action.
Inspired by Octo [octo_2023] and vision transformers [Dosovitskiy2020], we consider a randomly initialized Llama2 model of 12 layers with a latent dimension of 768, which we name Llama2-Base. In addition, multiple works have shown that multimodal inputs can be aligned to large-language models [liu2023llava, han2023imagebindllm, fu2024tvl, brohan2023rt2, mirchandani2023large]. Multi-modal language model, Palm-E [Driess2023] has shown success in enhancing generalization when being directly incorporated into robotic control [brohan2023rt2]. Therefore, we also investigate the effectiveness of using a large-language model for in-context robot learning by initializing the transformer with a pre-trained Llama2-7B. Due to the large domain gap between natural language and robot trajectories, a frozen language model may not be sufficient. Therefore, similar to prior work in multimodal alignment, we fine-tune the language model with LoRA [hu2021lora], with a rank of 32. Due to compute resource limitations, we are unable to fully fine-tune the model.
Loss Function To provide more supervision signals so that the model can better respond to the trajectory “prompt” we provide at test time, we reference works in training multi-turn conversation chatbots [chiang2023vicuna, liu2023llava], where they only compute loss on the response generated by the chatbot, instead of the prompt. Recall that in Section IV-A, we randomly sample the subsequence of the concatenated trajectories as the prompt trajectory. Analogously, we only compute action prediction with L1-loss for the actions after the prompt trajectories.
A key advantage of this framework is its use of the transformer’s sequential processing capability. Instead of reprocessing the entire sequence history for each model evaluation, as seen in previous works [octo_2023, kim2024openvlaopensourcevisionlanguageactionmodel, brohan2022rt, brohan2023rt2], the model employs a key-value (KV) caching mechanism, as discussed in [touvron2023llama]. This mechanism stores previous outputs, allowing the model to compute only the outputs for the new token. This approach significantly reduces computational overhead, lowering the complexity from quadratic to linear relative to the sequence length. Empirically, ICRT can inference at 39.6 Hz, allowing it to perform real-time close-loop control.
V Experiments
@@
In this section, we design an experimental setup to evaluate the in-context learning capabilities of the proposed models and compare them against several baselines. Instead of focusing on the difficulty of learning a specific task primitive, we design the experiments to assess the policy’s ability to accomplish novel tasks based on the provided prompt trajectories.
@@
Experiment Design We consider two action primitives: a pick-and-place primitive and a poking primitive. For each action primitive, we design six unseen tasks (as defined in Section III), with three tasks evaluating in-domain object generalization (selected from yellow cube, red cube, black cube, pink bowl, and blue bear and three objects unseen during training (selected from radish, blue sponge, grey dog, and black dog).
@@
Each task has five difficulty tiers. In the pick-and-place task, the model must identify the correct object to grasp and where to place it in a multi-object or multi-placement scenario. The tiers include: 1) no distractors, 2) one distractor object, 3) two distractors, 4) three distractors, and 5) one distractor placement position. In the poking task, the robot closes the gripper, pokes the object, lifts the end-effector, and opens the gripper, with tiers involving 0-4 distractor objects.
@@
The pick-and-place task is scored with 0.5 for a correct pick and 1 for a successful placement. In the poking task, failure is marked if the wrong object is poked. The model has 25 seconds (375 steps) for retries. Each difficulty level is attempted once, and we report the average success rate per task, along with the average success rate and standard deviation across the six tasks for each action primitive.
@@
Models The default ICRT is a randomly initialized Llama2-Base model pretrained on DROID and fully fine-tuned on ICRT-MT. We evaluate the impact of model initialization and training datasets by introducing the following three variants: 1) ICRT-Llama2, a pre-trained Llama2-7B language model fine-tuned on ICRT-MT with LoRA; 2) ICRT (DROID), a randomly initialized Llama2-Base model trained only on the DROID dataset; and 3) ICRT (MT), a randomly initialized Llama2-Base model trained only on the ICRT-MT dataset.
@@
We consider 3 baseline models. We train a goal-conditioned policy, where the goal observations are always prepended to the sequence, and each sequence is from one trajectory. This resembles the normal goal-conditioned imitation learning setup. Additionally, we finetune Octo [octo_2023], the state-of-the-art goal-image and language conditioned policy, and OpenVLA [kim2024openvlaopensourcevisionlanguageactionmodel], the state-of-the-art language conditioned multi-task imitation learning policy. Octo is fine-tuned using their official fine-tuning recipe. We incorporate action chunking into OpenVLA by asking it to predict the next 16 actions, which performs better than vanilla OpenVLA which predicts only the next step. Both of these methods are representative of robot policies that use next-token prediction objectives.
@@
Prompt Generation For each task, we collect 3 demonstrations (with zero, one distractor object, a distractor placement for pick-and-place, or two distractor objects for poking) as the prompt in total before running the experiment. Please refer to the website for a visual example. During testing, a random demonstration is drawn as a prompt to assess the model’s ability to generalize to different prompts. It’s important to note that the environment setup during policy rollout differs from the prompts’ setup, ensuring that the evaluation measures the model’s understanding of task-relevant information from the prompt, rather than simply copying actions from it.
Pick and Place Poke Average
Goal Condition 33.3 ($\pm$6.5) 6.7 ($\pm$4.6) 20.0 ($\pm$4.3)
Octo [octo_2023] 5.0 ($\pm$2.7) 13.3 ($\pm$6.2) 9.2 ($\pm$3.5)
OpenVLA [kim2024openvlaopensourcevisionlanguageactionmodel] 11.7 ($\pm$4.6) 3.3 ($\pm$3.3) 7.5 ($\pm$2.9)
ICRT 65.0 ($\pm$7.3) 93.3 ($\pm$4.6) 79.2 ($\pm$4.6)
TABLE I: Main Results. ICRT outperforms two state-of-the-art robot foundation models that are conditioned on goals or language in both pick-and-place and poking tasks. We evaluated each task primitive using six tasks not seen during training, conducting five trials per task for a total of 30 trials per primitive and 60 trials overall to calculate average performance. For each model, we report the mean success rate for each task, the overall success rate, and the corresponding standard error in parentheses.
@@
Results We present the results in Table I. For the pick-and-place primitive, we observe that the goal-conditioned policy generally succeeds in identifying the correct object to grasp when no distractor objects are present. However, its performance degrades significantly as the number of distractors increases. When the goal image only specifies the task but not the specific way to achieve it in the current environment, goal-conditioned policies often fail to execute the task effectively.
@@
The results suggest that ICRT outperforms the goal-conditioned policy in identifying the correct object to pick up and the appropriate placement location. The poking task presents a significant challenge for the goal-conditioned policies, as the goal position often closely resembles the start configuration. However, after conditioning on the prompt trajectory, ICRT is able to correctly identify the task as poking, and the results indicate that it consistently reaches the correct target object while ignoring distractors. Despite this, we do observe some failure modes with ICRT, such as missing the grasp of the target object, grasping the wrong object, or placing objects in incorrect locations. Specifically, when a distractor object shares the same color but has a different shape, the model struggles to accurately determine which object to grasp. This implies that additional fine-tuning of the vision encoder might be required to enhance model performance, a conclusion also reached by OpenVLA [kim2024openvlaopensourcevisionlanguageactionmodel].
Pick and Place Poke Average
ICRT-Llama2 43.3 ($\pm$7.9) 73.3 ($\pm$8.2) 58.3 ($\pm$6.0)
ICRT (DROID) 0.0 ($\pm$0.0) 0.0 ($\pm$0.0) 0.0 ($\pm$0.0)
ICRT (MT) 76.7 ($\pm$7.1) 70.0 ($\pm$8.5) 73.3 ($\pm$5.5)
ICRT +Prompt Loss 21.7 ($\pm$6.2) 23.3 ($\pm$7.9) 22.5 ($\pm$5.0)
ICRT 65.0 ($\pm$7.3) 93.3 ($\pm$4.6) 79.2 ($\pm$4.6)
TABLE II: Ablation Study. We ablate three key design choices: fine-tuning ICRT using a language model, training solely on the relevant dataset or the fine-tuning dataset, and the impact of including prompt loss in the training process. We report the mean success rates and their corresponding standard errors for the pick-and-place and poke tasks, as well as the overall average performance.
VI Ablations
@@
In this section, we provide additional experiments presented Table II that ablate on a few core design choices. We provide additional ablation studies on the official website.
VI-A Model Initialization
@@
We conducted ablation studies to examine the impact of using a pretrained Llama2 on language data and fine-tune it for robot sensorimotor sequence modeling. The results, presented in Table II, show that although ICRT-Llama2-7B achieves a lower training loss, its performance is worse compared to its smaller counterparts. This discrepancy may be attributed to a lower inference frequency of ICRT-Llama2 (10.7 Hz vs 39.6 Hz). Future work can focus on optimizing the inference speed of ICRT-Llama2 to improve performance.
VI-B Training Dataset
@@
We find that training on the DROID subset (see Section IV-A) is insufficient for completing any of the test tasks; the policy (ICRT (DROID)) shows no progress across all tasks. This suggests that although the DROID subset may offer greater visual diversity, the unique structure of ICRT-MT—where multiple tasks are performed from the same initial observation—is particularly beneficial in developing the in-context learning capabilities of a next-token prediction robot model.
@@
ICRT (MT) shows similar performance to ICRT that is pre-trained on DROID, especially for the pick-up and place primitive, even surpassing ICRT on the put radish in grey bowl task. However, ICRT (MT) does not perform as well on the poking primitive. The results suggest that it may be beneficial to pre-train the autoregressive model on a large dataset, as a diverse dataset may help the transformer to perform better alignment between visual features and control.
VI-C No Prompt Loss
@@
Following the design of many multi-turn conversation large language models or vision language model fine-tuning works [liu2023llava, chiang2023vicuna, liu2023improvedllava, dai2024instructblip], we do not calculate the loss for the predicted action in the prompt trajectories but only do so on the predictions after the prompt trajectories. We mark the model that calculates loss on the prompt as ICRT +Prompt Loss and the default model as ICRT. The results are shown in Table II. We find that only predicting the trajectories after the designated prompt trajectories can significantly improve the model’s performance. We hypothesize that in the situation where there is a loss on the prompt trajectories, the model is forced to do unconditional generation based on current observations for those prompts. This may cause the model to stop paying attention to the prompt, especially when there are multiple possible tasks available.
VII Limitations and Conclusion